Circular No. 9083

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 T3 (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to show cometary appearance by other CCD astrometrists. L. Buzzi, Varese, Italy, writes that images taken in moderate seeing with a 0.60-m reflector on Oct. 15.1 UT show a 12′′ coma that appears slightly elongated to the southwest, but with no clear tail. E. Guido, Castellammare di Stabia, Italy, reports that twenty co-added 120-s images taken remotely by G. Sostero and himself on Oct. 15.4 with a 0.25-m reflector near Mayhill, NM, U.S.A., show a coma nearly 12′′ in diameter, slightly elongated toward the southwest. W. H. Ryan (Magdalena Ridge Observatory 2.4-m reflector) reports a tail at p.a. ∼ 270^{o} on images taken during Oct. 15.44–15.45. R. Miles (2.0-m ‘Faulkes Telescope North’ at Haleakala, Oct. 15.5) reports a very diffuse coma of diameter 15′′, offset in p.a. ≈ 260^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Oct. 14.43092 5^{h}25^{m}15.^{s}46 +22^{o}02′04′′.2 19.3 LINEAR

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-T56.

T = 2010 Jan. 5.205 TT  ω = 29.^{o}066
Ω = 60.259 }2000.0
q = 2.33809 AU  i = 148.533

COMETS C/2009 H4–H8 (SOHO) Further to IAUC 9080, additional Kreutz sungrazers have been found

on SOHO website images. C/2009 H4 was small, slightly diffuse, and slightly elongated (peak mag ∼ 8). C/2009 H5 was diffuse with a short, faint tail (mag ∼ 7.5). C/2009 H6 (mag ∼ 7), C/2009 H7 (mag ∼ 8.5), and C/2009 H8 (mag ∼ 7–7.5) were all small and stellar in appearance.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 H4 Apr. 19.476 1^{h}56.^{m}9 +10^{o}16′ C2 MK 2009-N33
C/2009 H5  22.351 2 03.9 +10 50 C2 BZ 2009-N34
C/2009 H6  23.976 2 12.8 +11 43 C2 BZ 2009-N34
C/2009 H7  24.243 2 14.1 +11 46 C2 RK 2009-N34
C/2009 H8  24.471 2 22.5 +11 00 C3 RM 2009-N34
2009 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green